OpenStudy (anonymous):

simplify ln e ^x + 4y The answer options are 4xy, x/4y, e^x+4y, or x+4y.

OpenStudy (lgbasallote):

hint: \(\ln e = 1\)

OpenStudy (lgbasallote):

hint #2: use power rule on \(\ln e^x \) then let \(\ln e = 1\)

OpenStudy (anonymous):

The answer options are 4xy, x/4y, e^x+4y, or x+4y.